the roof of a tower is a square pyramid with side length 10 ft. The height of the pyramid is 6 ft. To the nearest square foot, find the area of the roofing material needed to cover the roof
Answer by lwsshak3(11628)

the roof of a tower is a square pyramid with side length 10 ft. The height of the pyramid is 6 ft. To the nearest square foot, find the area of the roofing material needed to cover the roof.
...
For this problem we will be working with two triangles in different planes.
The pyramid roof is made up of 4 identical triangles. Each triangle has a base=10 ft and a height=perpendicular line drawn from the apex of the pyramid to the base (in the same plane). Let's call this line h.
Looking from the side in a different plane we can see a right triangle with legs of 6 ft and 5 ft and a hypotenuse=h.
...
Calculations:
h=sqrt(6^2+5^2)=√(36+25)=√61
Area of 1 roof triangle=1/2*10*√61=39.05
Area of entire roof=39.05*4=156.2 sq ft